OpenStudy (anonymous):

In right triangle ABC in which c is the right angle prove sin2a=(2ab)/(c^2)

OpenStudy (asnaseer):

do you know how to expand \(\sin(2a)\)?

OpenStudy (anonymous):

yes I it 2sinacosa

OpenStudy (anonymous):

 \[\sin(2a) = 2\sin(a)\cos(a)\] SO you have to find sina and cosa..

OpenStudy (asnaseer):

correct - now just fill-in the blanks :)

OpenStudy (anonymous):

So would sin A be a/c?

OpenStudy (asnaseer):

yes

OpenStudy (anonymous):

and then the proof is complete?

OpenStudy (anonymous):

find cosa also..

OpenStudy (asnaseer):

almost - you also need cos(A)

OpenStudy (anonymous):

cosA would be b/c

OpenStudy (asnaseer):

yes

OpenStudy (anonymous):

just multiply sina cosa and 2..

OpenStudy (asnaseer):

now just plug those into the expansion of sin(2a)

OpenStudy (anonymous):

2 (a/c)(b/c)

OpenStudy (anonymous):

I get to 2(a/c)(b/c) plugging them in

OpenStudy (asnaseer):

thats it - well done! :)

OpenStudy (anonymous):

thanks for the help

OpenStudy (asnaseer):

yw :)